Fermi National Laboratory
 

Accelerator Update for November 3, 2006 - November 6, 2006

 

Friday November 3

The day shift began with the Tevatron (TeV) in access to complete the sector E4 vacuum repairs, with Operation monitoring the Antiproton Source (Pbar) stack, with Linac operating normally, with Booster in access, with the Recycler off due to a Main Injector (MI) access, and with MiniBooNE, NuMI, the Collider Detector at Fermilab (CDF), and D-Zero (D0) all in access.

Booster came out of access at 10:43 AM. Operations had beam established to the Booster dump by 11:18 AM.

At 1:27 PM, mechanical technicians reported a vacuum leak on the cryo system single-phase helium line in between a TeV sector E43 spool piece and a dipole magnet.

Booster experts began a study period at 3:10 PM.

Operators interlocked MI at 6:58 PM.

Saturday November 4

The midnight shift began with the TeV work continuing, with Operations monitoring a small stack (25.0E10), with Linac and Booster operating normally, and with all other beam lines and experiments off, in access, or waiting for beam.

At 2:30 AM, the I- Source tripped off for the first of three times in about an hour. It reset each time.

At 1:43 PM, a MI RF station (MIRF5) tripped off on a LCW flow fault and wouldn’t reset. Operators bypassed the station.

Operations began ramping MI at 2:10 PM.

Operations reported at 8:59 PM that the final vacuum certification was underway in TeV sector E43.

Operations prepared MI for access at 9:34 PM to inspect MIRF5.

A TeV expert completed a TeV hipot at 10:39 PM.

At 11:10 PM, RF experts reported that MIRF5 had an internal cavity LCW leak. Operations called in LCW technicians to investigate.

Sunday November 5

The midnight shift began with the TeV work continuing, with Operations monitoring a small stack, with Linac and Booster operating normally, and with all other beam lines and experiments off, in access, or waiting for beam.

Vacuum experts certified the vacuum at TeV sector E4 at 12:33 AM.

At 4:13 AM, Operations reported that the I- Source had suffered several high voltage and vacuum trips. They all reset okay.

MI experts reported that the MIRF5 cavity was flooded. The leak was fixed. Operations began the MI turn on.

Operations received permission to begin cooling down the TeV sector E4 cryo system at 9:39 AM.

Operations established beam to MiniBooNE at 11:34 AM. Unfortunately, the beam was intermittent because Booster was having a problem capturing beam from Linac.

Operation established beam to the Accumulator at 1 PM. Pbar began stacking at 1:56 PM.

At 1:58 PM, MI suffered the first of three vacuum bursts over an hour and twenty minutes. It reset after each trip.

At 2:33 PM, the 400MeV area Debuncher stopped regulating. Booster experts investigated and Operations contacted a Linac expert. A Linac tech arrived at 3:15 PM.

Operations halted the Linac beam at 3:25 PM to investigate the Debuncher problem.

Operations prepared MI at 4:43 PM for an access to investigate the vacuum bursts.

Linac experts reported at 6:09 PM that the Debuncher problems were due to LCW regulation problems.

Earlier, a NuMI kicker (KPS6) wouldn’t turn on. At 6:07 PM, Operations reported that a resister had burned up. The resistor was replaced but the kicker still wouldn’t turn on. The investigation continued.

Operations put MI into ramp mode at 6:52 PM.

Operations established beam to MiniBooNE and Pbar at 7:07 PM.

NuMI Kicker experts reported at 8:09 PM that they had found and fixed two bad connections.

NuMI Kicker experts turned on KPS6 at a lower that normal voltage at 8:35 PM and ran it that way for about forty-five minutes.

The Recycler completed a stash of antiprotons at 9:17 PM.

Kicker experts reported at 9:21 PM that they had completed their repairs on the NuMI kicker. Operations established beam to NuMI.

At 10:01 PM, Operations reported that the Booster RF system was sparking because there was still a problem with the Klystron Debuncher phase. Experts continued to work on the problem.

Monday November 6

The midnight shift began with the TeV sector E4 cryo system cooling down, with Linac having Debuncher problems, with the Booster and MI operating normally, with Operations monitoring a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:18 AM, Operations reported a Booster power supply (HS6) wasn’t regulating correctly. An expert was called in to investigate.

At 1:18 AM, Operations reported that the Pelletron was tripping off. Operations decided to leave it off for the night.

At 1:25 AM, a Linac expert reported that the Klystron Debuncher system had been out of phase for the last two days, which was causing the Booster problems.

Operations reported reestablishing beam to Pbar and NuMI at 1:42 AM.

The Recycler completed a stash at 3:50 AM.

The Recycler completed a stash at 7:03 AM.

At 7:34 AM, an Accumulator kicker (A:IKIK) tripped off and would not reset. Operators called an expert.

The Plans for Monday

The plans for today include getting the TeV electrical signoff completed and then beginning the TeV startup later this morning. D0 experimenters will continue working in their collision hall until late this afternoon. TeV experts will conduct TeV commissioning studies for 8 to 12 hours. The Run Coordinator would like the TeV to be in shot setup by late this evening and then establish a store on the Tuesday midnight shift.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/06/2006    email Fermilab